# Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads

### v4

@@ Abstract
Among ML operators today, GEneralMatrix Multiplication (GEMM)-based operators are known to be key operators that build the main backbone of ML models. As their computational overhead dominates the overall execution time (e.g., 42.8% - 96.6% in our results), GEMM operators have been the prime optimization targets for fast ML inference. This led to advanced GPUs and accelerators available today, which provided significant boost in the GEMM performance compared to CPUs, aligned with the lesson from Amdahl’s law. However, accelerating GEMM has significantly shifted the Amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is not dominant. Although the importance of non-GEMM performance is increasing, we have little knowledge about the non-GEMM performance horizon in the latest hardware platforms and models.
Therefore, to guide non-GEMM-oriented optimizations, we conduct a thorough performance analysis of 16 widely adopted ML models in Hugging Face and Torchvision on workstation and data center platforms with/without GPUs. We discover that non-GEMM performance bottleneck is a considerable issue across all the platforms and models, accounting for 11.3% to 73.6% of total latency, on average. The challenge significantly aggravates when we apply quantization, which is a common model compression technique, due to the boosted GEMM performance and extra non-GEMM operators for dequantization and requantization. To provide insights into non-GEMM optimization targets, we demystify the most dominant non-GEMM operators for each model and deployment software. We also show that widely adopted optimizations such as operator fusion do not completely address the non-GEMM performance bottleneck, where non-GEMM operators still account for 15% to 48% of total latency. We will open-source our non-GEMM-oriented benchmark framework to facilitate research in non-GEMM optimization.
@@ I Introduction
The success of machine learning (ML) in various problem domains, such as computer vision (CV)  [22, 21, 54, 32, 37] and natural language processing (NLP) [31, 4, 58], made ML workloads pervasive in various computing platforms from edge to cloud devices. ML model inference involves billions of multiply-and-accumulate (MAC) operations (e.g., 497 billions of MAC operations for ResNet 50 [22]). Such MAC operations originate from GEneral Matrix Multiplication (GEMM)-based operators, such as CONV2D, Linear, and BMM (batched matrix multiplication). The GEMM-based operators dominate in terms of the total execution time on CPUs, as shown in Figure 1. Therefore, GPUs and accelerators have focused on the optimization of the GEMM-based operators, which significantly enhanced the computational performance (e.g., latency and throughput) of end-to-end ML model inference, as highlighted in Figure 1.
@@
However, because the GEMM operators are being accelerated, the non-GEMM operators, such as memory operations (e.g., reshape, view, and transpose), normalization, and logit computation functions (e.g., Softmax), now account for a considerable amount of the end-to-end latency, compared to that of GEMM operators.  Figure 1 shows the profiled latency breakdown into GEMM and non-GEMM operators running inferences on state-of-the-art large language (GPT2-XL [4]) and image classification (Swin Transformer [37]) models. The motivational data imply a major shift of the landscape of Amdahl’s law [24] for ML inference acceleration, indicating that we now need to consider non-GEMM operators as one of the major optimization targets in the ML system optimization. However, the research community today lacks a thorough and systematic performance analysis and characterization of non-GEMM operators in the latest models, which hinders the development of non-GEMM oriented optimization techniques.
Therefore, we collect widely-used ML models from Hugging Face [61] and Torchvision [38] and perform a thorough performance characterization of of non-GEMM operators in the 18 latest models of four major task domains: Image Classification (IC), Image Segmentation (IS), Object Detection (OD), and Natural Language Processing (NLP). We evaluate the effect of GPU acceleration on the relative latency across GEMM and non-GEMM operators in collected models and conduct deep-dive analysis on the impact of different hardware platforms (mobile, workstation, and datacenter), deployment software, and common optimizations (operator fusion and quantization). Based on our case studies, we highlight that the non-GEMM performance challenge is common in accelerated inferences and existing optimization techniques (e.g., operator fusion) cannot completely address the challenges. Also, we demystify the most time-consuming non-GEMM operators in each model, which will help the research community identify non-GEMM operators to be optimized.
To facilitate such research in non-GEMM-oriented optimization techniques, we build an open-source benchmark specialized in non-GEMM performance anlysis, NonGEMM Bench, which will be released after publication. NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch [46], ONNX [7], and TensorRT [43], in addition to the preset of non-GEMM operators collected from the selected 16 popular models, which provides desired flexibility to users for follow-up research.
@@
  • We evaluate the performance of non-GEMM operators with quantization and show the non-GEMM bottleneck aggravates with data.
  • We will open-source NonGEMM Bench, an expandible benchmark flow that enables thorough non-GEMM performance characterization for any model supported by ONNX runtime, Tensor RT, and PyTorch, to facilitate non-GEMM-oriented research.
@@ II Background
Figure 2: A description of Linear (a) and Conv1D (c) operators as GEMM operators example, and that of non-maximum suppression [21] (b) and Layer Normalization [6] (d) as a non-GEMM operator example.
@@ II-A ML Operators
ML operators are the building blocks for ML models, which define the computation over one or multiple input tensors. Examples include convolution (Conv2d), matrix multiplication (linear, BMM, etc.), activation, and normalization, as listed in Figure 2. We categorize operators into two classes: GEMM operation-based and the others, which we term as non-GEMM operators. We discuss each class of ML operators next.
II-A1 GEMM-based Operators (GEMM Operators)
@@
GEMM-based operators (or GEMM operators) refer to all the ML operators that can be represented as a matrix multiplication operation, which include linear, Conv2d, and batched-matrix multiplication (BMM). For example,  Figure 2 (a) and (c) illustrate two popular GEMM operators: Linear and Conv2d operators, respectively. Each operator can be represented into a perfectly nested loop with multiply-and-accumulate (MAC) operation in the inner-most loop. Note that variants that are not matrix multiplication in the default form like Conv2d can be converted into GEMM (e.g., im2col [11]), which motivated the term, GEMM operator.
@@
GEMM-based operators are known to be compute-intensive, which accounts for the majority of the execution unless accelerated by GPUs or accelerators, as CPU results in in Figure 1 show. However, they have regular computation patterns that can be summarized as a perfectly nested for loop. The regular pattern allows various loop optimization techniques such as loop reordering, tiling, and parallelization, which is referred to as dataflow [45, 33, 66] With the dominance of GEMM operators in execution time and high optimization potential together, GEMM operators have been the prime optimization target for acceleration, which led to high-performance GPUs (e.g., H100 [13]) and accelerators [29].
II-A2 Non-GEMM Operators
@@
Non-GEMM operators refer to all ML operators other than GEMM operators. They span various functionalities (e.g., memory layout manipulation and normalization) other than applying weights to input tensors. Because of their diverse functionalities, their computation patterns are often not a perfectly nested loop with MAC, which can also involve non-linear functions and memory-oriented operations. For example,  Figure 2 (b) shows non-maximum suppression (NMS) operator often found in R-CNN model variants [21, 53]. As found in the example, the entire operation cannot be summarized into single perfectly-nested loop, which involve other operations such as sort and filtering. In addition, the operation involves a conditional statement, which introduces non-deterministic behaviors to the operator. The layer normalization example in Figure 2 (b) also shows another key characteristic of the non-GEMM operators: non-linear functions. Because of such characteristics distinguished from GEMM operators, optimization methodologies for GEMM operators cannot be applied to accelerate non-GEMM operators.
@@
To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 16 recent models in the computer vision and natural language processing domains. We select models based on their popularity in the Hugging Face to obtain realistic workload. We list the models we investigated in Table I. Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table II.
@@
  • Activation. Activation operators introduce non-linearity into the model. Rectified Linear Unit (ReLU) function [41] is an example of activation operators used in CNN based ML models  [54, 22, 55]. It injects non-linearity into the model based on the sign of the data by applying the function $X_{i}=Max(0,X_{i})$ element-wise to the data, where $X_{i}$ is the i-th data point. Another variant of activation operators is the the Gaussian Error Linear Units function (GELU) [23] which is a popular activation function adopted in transformer based ML models s [4, 64, 32, 37]. Unlike ReLU, GELU accounts for the value of the data when inserting the non-linearity and not only the sign. It multiplies the input $X_{i}$ by the Cumulative Distribution Function (CDF) of a Gaussian distribution: $GELU(X_{i})=X_{i}*\phi(X_{i})$ [23] where $X_{i}$ is the i-th data and $\phi$ is the CDF of the Gaussian distribution.
@@
  • Element-Wise Arithmetic. Elemennt-wise arithmetic operators cover non-GEMM operators applied on individual elements in a tensor. For example, Figure 3 (c) contains an element-wise division applied to scale the elements of tensors in the attention block
  • RoI Selection. RoI selection operators are found in R-CNN variant detection models. [53, 21]. They filter down bounding boxes proposed by the region proposal network (Figure 3 (b)) and align the remaining boxes to the objects detected in the image. Non-Maximum Suppression (NMS) is an example of RoI Selection, which is described in Figure 2. Given a list of scores and bounding box information, it selects bounding boxes by applying the Intersection over Union (IoU) metric.
@@
II-B ML Models and Popular Tasks
The heterogeneity in non-GEMM operators enabled ML developers to build models supporting a wide range of tasks (e.g. Computer Vision and NLP). As highlighted in  Figure 3, computer vision ( Figure 3 a and b) and NLP ( Figure 3) models are characterized by distinct architectures leveraging unique combinations of GEMM and non-GEMM operators.
For example, classical models in image classification shown in  Figure 3 (a) consist of deep convolution networks (CNN). which cascade Conv2d (GEMM) with normalization and activation operators [54, 22]. Object Detection models also rely on convolution network for feature extraction, region proposal and classification. However, they pair the CNNs with ROI selection non-GEMM operators to process and filter the detections (Figure 3 (b)). On the other hand, NLP and language models rely on the transformer architecture, which leverages the attention mechanism introduced in [59]. Transformers combine BMMs and Linear GEMM operators with normalization, memory and arithmetic non-GEMM operators as shown in Figure 3. This diversity in non-GEMM operators and model architectures across tasks paired the preliminary data presented in  Figure 1 motivate us to study the impact of non-GEMM operators have on the inference performance of popular models.
Figure 4: An overview of NonGEMM Bench flows
III Performance Characterization Methodology
To understand the realistic performance landscape of the latest ML models with non-GEMM workloads, we must (1) capture operator level performances in end-to-end inferences to understand, (2) use widely-used models by the research community and industry, (3) cover diverse task domains, and (4) use real datasets.
However, ML Benchmarks available today (e.g., MLPerf [52]), unfortunately, do not satisfy all the requirements since they do not focus on the non-GEMM operators. Long-tail bench [34] identified a similar problem as this work, but it focuses on a limited set of custom kernels, which fails to represent broad task domain. Therefore, to facilitate our non-GEMM operators analysis and better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench. NonGEMM Bench provides operator-level breakdown of end-to-end inference latency in the operator graph level, which enables detailed non-GEMM operator performance analysis, as we present in Section IV. To capture the performance in the latest ML workload, we select 15 highly downloaded (more than 10K downloads on average) models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench and our analysis. We discuss the datasets and models in NonGEMM Bench in detail and describe the structure of NonGEMM Bench next.
III-A Evaluated Models
Table I lists the NonGEMM Bench model registry which contains 15 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 0.85M to 7B, demonstrating the diverse model coverage of NonGEMM Bench. The selected models cover four major task domains in ML, which include Image Classification (IC), Object Detection (OD), Image Segmentation (IS), and Natural Language Processing (NLP).
III-A1 Computer Vision (CV)
Deep learning is widely used in the CV domain [32]  [22]  [53]. NonGEMM Bench covers three important tasks in the domain:
Image Classification (IC). Image classification refers to a CV task that identifies a class label of a given image. Image classification models extract features (i.e., high-level and dimensional information of the input image) from an input image and report the class label utilizing the features. NonGEMM Bench includes eight most popular IC models in HuggingFace [1]: Three variants of Vision Transformer [32], and three variants of Swin Transformer  [37].
Image Segmentation (IS). Image segmentation refers to a computer vision task that identifies the area in a image for each class. Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels that belong to each object. NonGEMM Bench includes two state-of-the-art IS models: Segformer  [64] and MaskFormer  [10]. Object Detection (OD). Object detection refers to a computer vision task that identifies the location of objects in an image and outputs the bounding box of each object. OD models extract features and generate region proposals, which refer to the candidate locations and bounding boxes of objects. Using Region of Interest (RoI) processing non-GEMM operators in Table II, OD models refine raw region proposals generated by a region proposal network. The refined RoIs are used as inputs to the CNN classifier at the end, and the classifier determines the class label of objects inside each refined RoI. NonGEMM Bench includes three popular OD models [2]: FasterRCNN  [53], MaskRCNN  [21], and DETR  [9].
III-A2 Natural Language Processing (NLP)
NLP refers to tasks involving the analysis and understanding of human (natural) language. NLP models extract context and features from an input text sequence and use the extracted context and features to perform multiple applications like translation and text generation [68] [8]. Transformer [59] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state-of-the-art generative large language models like GPT  [50] and Llama  [58].  Figure 3 (c) shows the layer architecture of GPT’s transformer. It consists of a self-attention block built by cascading GEMM operators with Normalization, Memory, Logit Comuputation and Element-wise Arithmetic non-GEMM oeprators ( Table II). NonGEMM Bench includes five popular NLP models  [3]: Bert [31], three variants of GPT2 [4], and Llama2-7b [58].
III-B Profiling Methodology
To enable the thorough analysis of the landscape of non-GEMM performance presented in Section IV, NonGEMM Bench provides an end-to-end profiling flow to breakdown the inference latency at the operator level. We discuss input, output, and the benchmark software structure.
III-B1 NonGEMM Bench Inputs
NonGEMM Bench receives target deployment software flow (PyTorch, ONNX runtime, or TensorRT), dataset, and other configurations listed in Figure 4 as inputs. Optionally, users can provide custom models with associated datasets to profile their workloads. By default, NonGEMM Bench runs the 15 models in Table I. We describe each input in detail as follows.
Models. As described in  Section III-A, NonGEMM Bench includes a registry of 15 selected popular ML models. Nonetheless, we designed NonGEMM Bench to be easily expandable to accommodate rapidly evolving ML models that constantly introduce new operators. Users can benefit from the features of our benchmark by simply plugging their new models into the NonGEMM Bench model registry (Figure 4) by specifying the model class, its weights and any data preprocessor.
Deployment Flow. NonGEMM Bench deploys its models using three popular inference frameworks: ONNX Runtime [17], PyTorch [46] TensorRT [43].
Datasets. To evaluate the models, NonGEMM Bench utilizes real data from popular datasets in each domain. We use ImageNet 2012  [15] and MS COCO  [36] for computer vision tasks. As for language models, we use wikitext dataset [39] available on HuggingFace. For custom models, users can specify their own dataset for the model.
Misc. Configurations. This includes the batch size, the number of profiling iterations, and the target device.
TABLE I: Tasks and Models Evaluated in Section IV
Application Models # Params Dataset
Image Classification (IC) ViT base (Vt-b) [32] 307M ImageNet  [15]
ViT large (Vt-l) [32] 307M
ViT huge (Vt-h) [32] 632M
Swin tiny (Sw-t) [37] 29M
Swin small (Sw-s) [37] 50M
Swin base (Sw-b) [37] 88M
Object Detection (OD) FasterRCNN (FRCNN) [53] 42M COCO  [36]
MaskRCNN (MRCNN) [21] 44M
DETR [9] 41M
Image Segmentation (IS) Maskformer (MF) [10] 102M COCO  [36]
SegFormer (Seg) [64] 3.7M
Natural Language Processing (NLP) GPT2 [4] 117M wikitext  [39]
GPT2 Large (gpt2-l) [4] 762M
GPT2 X-Large (gpt2-xl) [4] 1.5B
Llama 2-7B [58] 7B
Bert [31] 110M
III-B2 End-to-End Inference Flow
The end-to-end inference flow profiles NonGEMM Bench models running inferences on data samples from the selected dataset.
First, the Graph Extractor extracts the operator graph of a given model based on the selected deployment flow. NonGEMM Bench calls graph exporters from the HuggingFace transformers library [62] or from the PyTorch libraries.
Second, model specific Data Preprocessing functions fetch the raw data from the specified dataset, clean the data, and apply tokenization or image to tensor transformations to generate input tensors for the profiling stage.
Third, End-to-End Inference Profiling functions run the profiling iterations on the received the model and its inputs. The profiling functions are deployment flow specific. For ONNX Runtime, we invoke Execution Provider profiling in the inference session. Currently, we only support operator level profiling for CPU and CUDA execution providers [17] because profiling support to other ONNX RUNTIME backends is not stable yet. For PyTorch, we use the PyTorch Profiler [49]. For TensorRT, we leverage the TensorRT open source software framework profiling APIs [44].
III-B3 NonGEMM Bench Outputs
NonGEMM Bench generates many statistics organized into three categories: performance, workload, and non-GEMM-specific reports. The performance report includes key performance metrics such as the end-to-end latency with operator level break-downs and peak memory usage. The workload report includes the types of operators and the shape of the tensors for each operator captured during inference on realistic data. Finally, the non-GEMM report provides insights on non-GEMM operators, such as the number of operator variants of the same class of non-GEMM operator (e.g., layernorm implementation could have multiple variants for each ML framework) and non-GEMM operator trace on different domains.
@@ III-B3 NonGEMM Bench Outputs
TABLE II: Non-GEMM operators in eight selected model variants from Table I and their characteristics.

### v5

@@ Abstract
Among ML operators today, GEneralMatrix Multiplication (GEMM)-based operators are known to be key operators that build the main backbone of ML models. As their computational overhead dominates the overall execution time (e.g., 42.8% - 96.6% in our results), GEMM operators have been the prime optimization targets for fast ML inference. This led to advanced GPUs and accelerators available today, which provided significant boost in the GEMM performance compared to CPUs, aligned with the lesson from Amdahl’s law. However, accelerating GEMM has significantly shifted the Amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is now significant. Although the importance of non-GEMM performance is increasing, we have little knowledge about the non-GEMM performance horizon in the latest hardware platforms and models.
Therefore, to guide non-GEMM-oriented optimizations, we conduct a thorough performance analysis of 17 widely adopted ML models in Hugging Face and Torchvision on workstation and data center platforms with/without GPUs. We discover that non-GEMM performance bottleneck is a considerable issue across all the platforms and models, accounting for 11.3% to 73.6% of total latency, on average. The challenge significantly aggravates when we apply quantization, which is a common model compression technique, due to the boosted GEMM performance and extra non-GEMM operators for dequantization and requantization. To provide insights into non-GEMM optimization targets, we demystify the most dominant non-GEMM operators for each model and deployment software. We also show that widely adopted optimizations such as operator fusion do not completely address the non-GEMM performance bottleneck, where non-GEMM operators still account for 15% to 48% of total latency. We will open-source our non-GEMM-oriented benchmark framework to facilitate research in non-GEMM optimization.
@@ I Introduction
The success of machine learning (ML) in various problem domains, such as computer vision (CV)  [25, 24, 51, 33, 38] and natural language processing (NLP) [19, 7, 56], made ML workloads pervasive in various computing platforms from edge to cloud devices. ML model inference involves billions of multiply-and-accumulate (MAC) operations (e.g., 497 billions of MAC operations for ResNet 50 [25]). Such MAC operations originate from GEneral Matrix Multiplication (GEMM)-based operators, such as CONV2D, Linear, and BMM (batched matrix multiplication). The GEMM-based operators dominate in terms of the total execution time on CPUs, as shown in Figure 1. Therefore, GPUs and accelerators have focused on the optimization of the GEMM-based operators, which significantly enhanced the computational performance (e.g., latency and throughput) of end-to-end ML model inference.
@@
However, because the GEMM operators are being accelerated, the non-GEMM operators, such as memory operations (e.g., reshape, view, and transpose), normalization, and logit computation functions (e.g., Softmax), now account for a considerable amount of the end-to-end latency, compared to that of GEMM operators.  Figure 1 shows the profiled latency breakdown into GEMM and non-GEMM operators running inferences on state-of-the-art large language (GPT2-XL [7]) and image classification (Swin Transformer [38]) models. The motivational data show that the non-GEMM operations now can account for the majority of the latency with GEMM acceleration, indicating that we now need to consider non-GEMM operators as one of the major optimization targets in the ML system optimization. However, the research community today lacks a thorough and systematic performance analysis and characterization of non-GEMM operators in the latest models, which hinders the development of non-GEMM oriented optimization techniques.
Therefore, we collect widely-used ML models from Hugging Face [59] and Torchvision [55] and perform a thorough performance characterization of of non-GEMM operators in the 17 latest models of four major task domains: Image Classification (IC), Image Segmentation (IS), Object Detection (OD), and Natural Language Processing (NLP). We evaluate the effect of GPU acceleration on the relative latency across GEMM and non-GEMM operators in collected models and conduct deep-dive analysis on the impact of different hardware platforms (workstation and data center), deployment software, and common optimizations (operator fusion and quantization). Based on our case studies, we highlight that the non-GEMM performance challenge is common in accelerated inferences and existing optimization techniques (e.g., operator fusion) cannot completely address the challenges. Also, we demystify the most time-consuming non-GEMM operators in each model, which will help the research community identify non-GEMM operators to be optimized.
To facilitate such research in non-GEMM-oriented optimization techniques, we build an open-source benchmark specialized in non-GEMM performance anlysis, NonGEMM Bench, which will be released after publication. NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch [44], ONNX [10], and TensorRT [4], in addition to the preset of non-GEMM operators collected from the selected 17 popular models, which provides desired flexibility to users for follow-up research.
@@
  • We evaluate the performance of non-GEMM operators with quantization and quantitatively show the non-GEMM bottleneck aggravates with quantization.
  • We open-source NonGEMM Bench, an extensible benchmark flow that enables thorough non-GEMM performance characterization for any model supported by ONNX runtime, TensorRT, and PyTorch, to facilitate non-GEMM-oriented research.
@@ II Background
Figure 2: Descriptions of example non-GEMM and GEMM operators. (a) (Non-GEMM) non-maximum suppression [24], (b) (GEMM) Conv1D (c) (Non-GEMM) Layer Normalization [9], and (d) (GEMM) Linear.
@@ II-A ML Operators
ML operators are the building blocks of ML models, which define the computation over one or multiple input tensors. Examples include convolution (Conv2d), matrix multiplication (linear, BMM, etc.), activation, and normalization, as listed in Figure 2. We categorize operators into two classes: GEMM operation-based (”GEMM operators”) and the others (”Non-GEMM operators”). We discuss each class of ML operators next.
@@
GEMM-based Operators (GEMM Operators). GEMM-based operators (or GEMM operators) refer to all the ML operators that can be represented as a matrix multiplication operation, which include linear, Conv2d, and batched-matrix multiplication (BMM). For example,  Figure 2 (b) and (d) illustrate two popular GEMM operators: Linear and Conv2d operators, respectively. Each operator can be represented into a perfectly nested loop with multiply-and-accumulate (MAC) operation in the inner-most loop. Note that variants that are not matrix multiplication in the default form like Conv2d can be converted into GEMM (e.g., im2col [14]), which motivated the term, GEMM operator.
@@
GEMM-based operators are known to be compute-intensive, which accounts for the majority of the execution unless accelerated by GPUs or accelerators, as CPU results in Figure 1 show. However, they have regular computation patterns that can be summarized as a perfectly nested for loop. The regular pattern allows various loop optimization techniques such as loop reordering, tiling, and parallelization, which is referred to as dataflow [43, 34, 62] With the dominance of GEMM operators in execution time and high optimization potential together, GEMM operators have been the prime optimization target for acceleration, which led to high-performance GPUs (e.g., H100 [48]) and accelerators [31].
@@
Non-GEMM Operators. Non-GEMM operators refer to all ML operators other than GEMM operators. They span various functionalities (e.g., memory layout manipulation and normalization) other than applying weights to input tensors. Because of their diverse functionalities, their computation patterns are often not a perfectly nested loop with MAC, which can also involve non-linear functions and memory-oriented operations. For example,  Figure 2 (a) shows non-maximum suppression (NMS) operator often found in R-CNN model variants [24, 50]. As found in the example, the entire operation cannot be summarized into single perfectly-nested loop, which involve other operations such as sort and filtering. In addition, the operation involves a conditional statement, which introduces non-deterministic behaviors to the operator. The layer normalization example in Figure 2 (c) also shows another key characteristic of the non-GEMM operators: non-linear functions. Because of such characteristics distinguished from GEMM operators, optimization methodologies for GEMM operators cannot be applied to accelerate non-GEMM operators.
@@
To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 17 recent models in the computer vision and natural language processing domains. We select models based on their popularity in the Hugging Face to obtain realistic workload. We list the models we investigated in Table II. Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table I.
@@
  • Activation. Activation operators introduce non-linearity into the model. Rectified Linear Unit (ReLU) function [41] is an example of activation operators widely used in CNN based ML models  [51, 25, 52]. ReLU injects non-linearity into the model based on the sign of the data by applying the element-wise function, $ReLU(X)=Max(0,X)$. Another variant of activation operators is the the Gaussian Error Linear Units function (GELU) [26], which is a popular activation function adopted in transformer based ML models [7, 61, 33, 38]. Unlike ReLU simply gates out negative values to be 0, GELU requires to compute the Cumulative Distribution Function (CDF) of a Gaussian distribution, which is often denoted as $\phi$. GELU multiplies the input $X$ by the Cumulative Distribution Function (CDF) of a Gaussian distribution ($\phi$): $GELU(X)=X*\phi(X)$ [26].
@@
  • Element-Wise Arithmetic. Element-wise arithmetic operators refer to all the operations applied on individual elements in a tensor (other than activations). For example, Figure 3 (c) contains an element-wise division applied to scale the elements of tensors in the attention block.
  • RoI Selection. RoI selection operators are found in R-CNN variants. [50, 24]. They filter down bounding boxes proposed by the region proposal network (Figure 3 (b)) and align the remaining boxes to the objects detected in the image. Non-Maximum Suppression (NMS) is an example of RoI Selection, which is described in Figure 2. Given a list of scores and bounding box information, it selects bounding boxes by applying the Intersection over Union (IoU) metric.
@@
TABLE I: Non-GEMM operators in eight selected model variants from Table II and their characteristics. Example input shapes are captured based on inferences using real datasets.